Y=5x-7
-3x-2y=-12

Answer by alicealc(293)

y = 5x - 7

substitute y = 5x - 7 into this equation:
-3x - 2y = -12
-3x - 2*(5x - 7) = -12
-3x - 10x + 14 = -12
14 + 12 = 3x + 10x
26 = 13x
x = 26/13 = 2

y = 5x - 7 = 5*2 - 7 = 10 - 7 = 3

so, the solution is x = 2, y = 3